# DynVFX: Augmenting Real Videos with Dynamic Content

### v1

@@ Abstract.
We present a method for augmenting real-world videos with newly generated dynamic content. Given an input video and a simple user-provided text instruction describing the desired content, our method synthesizes dynamic objects or complex scene effects that naturally interact with the existing scene over time. The position, appearance, and motion of the new content are seamlessly integrated into the original footage while accounting for camera motion, occlusions, and interactions with other dynamic objects in the scene, resulting in a cohesive and realistic output video. We achieve this via a zero-shot, training-free framework that harnesses a pre-trained text-to-video diffusion transformer to synthesize the new content and a pre-trained Vision Language Model to envision the augmented scene in detail. Specifically, we introduce a novel inference-based method that manipulates features within the attention mechanism, enabling accurate localization and seamless integration of the new content while preserving the integrity of the original scene. Our method is fully automated, requiring only a simple user instruction. We demonstrate its effectiveness on a wide range of edits applied to real-world videos, encompassing diverse objects and scenarios involving both camera and object motion [^1] Code will be made publicly available.. Project page: https://dynvfx.github.io/
@@
journal: TOG
Figure 1. DynVFX augments real-world videos with new dynamic content described via simple user-provided text instruction.
@@ 1. Introduction
Incorporating computer-generated imagery (CGI) into real-world footage has been a transformative capability in film production, enabling the creation of visual effects that would be difficult or impossible to achieve otherwise. For instance, the seamless integration of CGI characters, such as Gollum in The Lord of the Rings or T-Rex in Jurassic Park, has empowered filmmakers to blend fantastical elements with real-world environments, resulting in immersive storytelling. Inspired by these capabilities, we pose a new creative task: augmenting real-world videos with newly generated dynamic content. Specifically, given an input video and a text prompt describing the desired edit, our goal is to synthesize new dynamic objects or complex scene effects, which naturally interact with the existing scene across time (Fig. 1).
@@
Our task poses several new fundamental challenges. First, the generation must be content-aware, such that the position, appearance, and motion of the synthesized dynamic content integrate naturally with the original scene. This entails synthesizing objects that respect occlusions, maintain appropriate relative size and perspective with respect to the camera position, and realistically interact with other dynamic objects. All of this must be achieved while maintaining the integrity of the original video, ensuring that new content enhances the scene without compromising its authenticity.
Our method leverages a pre-trained text-to-video model without any fine-tuning or additional training. Specifically, given an input video along with a short user instruction describing the new content (e.g., “add a massive whale”), our method produces an edited video where the new content is seamlessly integrated into the original video. We achieve it by estimating the residual to the latent representation of the original video. To envision the edited scene and to identify prominent foreground objects and visual elements, we leverage a vision-language model that translates the user’s instructions into detailed prompts for the text-to-video diffusion model.
As our task requires careful placement of the new content, we propose to steer the localization of the edit through Anchor Extended Attention - incorporate a specific set of keys/values extracted from the original video as additional context to the model. Additionally, to improve the edit harmonization with the original scene, we propose to iteratively update the estimated edit residual.
@@
It is worth noting that our method utilizes a publicly available text-to-video model, which exhibits a significant gap in video generation quality compared to recent state-of-the-art video models. Nevertheless, we observe that within our problem formulation and objective, we can distill from this model surprisingly powerful generative capabilities. We demonstrate the effectiveness of our approach on a variety of edits applied to real-world videos.
@@
  • We introduce a new task of integrating newly generated dynamic content into real-world videos without relying on the user to provide complex references of the effect (for example, a VFX asset or masks to specify where to locate the VFX).
  • We propose a tuning-free, zero-shot method that enables harmonized content integration while maintaining high fidelity to the original scene
  • We propose an automatic VLM-based evaluation metric tailored for our task, considering multiple factors, including original content preservation, new content harmonization, overall visual quality, and alignment with the edit prompt.
  • We demonstrate state-of-the-art results compared to competing methods, achieving the best trade-off between synthesizing new dynamic elements and maintaining high fidelity to the original content.
@@ 2. Related Work
Text-to-Video Models. With the rise of large-scale video-text datasets, there has been significant progress in training text-to-video models using new architectures (Bar-Tal et al., 2024; Polyak et al., 2024). While the foundational architecture of diffusion models has been commonly linked to inflating text-to-image U-Net-based models with temporal layers (Wang et al., 2023a; cerspense, 2023; Guo et al., 2023), recently, a new family of Transformer-based models (OpenAI, 2024; Kong et al., 2025; Yang et al., 2024; Zheng et al., 2024; HaCohen et al., 2024), referred to as Diffusion Transformers (DiTs) (Peebles and Xie, 2023), have gained popularity, as DiTs enhance spatial coherence and enable arbitrary aspect ratio and video-length training. In this work, we utilize a publicly available DiT-based model (Yang et al., 2024), CogVideoX, for augmenting real-world videos with newly generated dynamic content in a zero-shot manner.
Figure 2. Controlling fidelity to the original scene using different extended attention mechanisms. (a-b) SDEdit suffers from the original scene preservation/edit fidelity trade-off. (c-e) Three Extended Attention variants during sampling demonstrate different control levels: Full Extended Attention closely reconstructs the input scene, Masked Extended Attention proves too constrained in overlapping regions despite allowing new content emergence, and our Anchor Extended Attn. achieves optimal results by applying dropout – extending attention only at sparse points within selected regions.
@@
Object Insertion in Images. With the advancement of text-to-image models, techniques for image manipulation leveraging such models have also evolved rapidly. Among these advancements, notable progress has been made in the task of instruction-based image editing. Several works (Zhang et al., 2023c; Sheynin et al., 2023; Brooks et al., 2023; Zhang et al., 2023a) have proposed to directly fine-tune generative models on pairs of original and edited images coupled with user-provided instructions. EmuEdit (Sheynin et al., 2023) leverages a diffusion model trained on a large synthetic dataset to perform various editing tasks guided by task embeddings.
@@
The task of object insertion into images belongs to the same category and can be considered a subfield of instruction-based editing methods. For instance, EraseDraw (Canberk et al., 2024), Paint By Inpaint (Wasserman et al., 2024), and ObjectDrop (Winter et al., 2024) leverage inpainting models to create paired-image datasets, which are then used to fine-tune image editing models. However, extending these approaches to videos presents significant challenges. In particular, generating large-scale instruction-paired video datasets can be prohibitively expensive both in time and computational resources, as it requires substantial manual effort to annotate frames and ensure alignment between textual instructions and video content. This cost and complexity make it challenging to adapt existing image-based methodologies directly to the video domain.
Concurrently to our method, Add-It (Tewel et al., 2024) proposes to manipulate the attention features of a pre-trained text-to-image diffusion model to insert objects into images in a training-free manner. While their method relies on weighted global extended attention, we propose to apply extended attention only to specific regions of the source scene, allowing the generation to focus on essential elements.
@@
Controllable Video Generation. Recently, numerous methods have been developed to incorporate various forms of control signals into video generation pipelines. Several video-to-video methods propose to condition the generation on per-frame spatial maps such as depth maps and edge maps (Wang et al., 2023b; Chen et al., 2023). A line of work (Yatim et al., 2023; Zhao et al., 2023; Park et al., 2024; Jeong et al., 2023; Geyer et al., 2023) proposed to utilize a text-to-video model for the task of motion transfer. Unlike these methods, which are not designed to deviate from the existing structures within a video, our approach focuses on integrating additional dynamic elements into the video.
Recent methods (Bar-Tal et al., 2024; Ma et al., 2024; Zhang et al., 2023b; Zi et al., 2024; Zhou et al., 2023) have explored adapting text-to-video models for video inpainting by conditioning on a masked video and a corresponding binary mask. This setup encourages the model to preserve unmasked information while generating new content in the masked region. All of these methods require user-provided masks — an impractical requirement for integrating complex dynamics as it requires anticipating the placement of dynamic objects (e.g., Fig. 5 jellyfish, tsunami, dinosaurs) ahead of time. In contrast, our method allows for automatic new content localization without any user-provided masks. While a static object can be masked with a simple bounding box, manually defining masks for complex motion or interactions per frame is extremely difficult.
Recently, generative Omnimatte (Lee et al., 2024) proposed a method to automatically decompose a video into object layers and their corresponding effects. However, it is not designed for the task of new content generation, as it allows only for the removal of existing scene elements. VideoDoodles (Yu et al., 2023) combines hand-drawn animations with video footage in a scene-aware manner by tracking a user-provided planar canvas in 3D. However, it does not support the creation of non-planar animations. Our framework includes both localizing dynamic new content and integrating in a scene-aware manner without reliance on user-provided positions.
@@
Language Models for Video Content Creation. Advancements in Vision Language Models (VLMs) have enabled methods to utilize such models in various video-related tasks. Some methods (Chen et al., 2024; Yang et al., 2024) use VLMs to produce detailed video captions from a series of frames, which are then utilized to train text-to-video generative models. Other methods utilize such models for achieving better generation controllability. For instance, VideoDirectorGPT (Lin et al., 2024) utilizes a VLM for multi-scene video generation by training diffusion adapters to incorporate additional conditioning inputs, while LVD (Lian et al., 2023) incorporates layout guidance from the VLM during the sampling process. AutoVFX (Hsu et al., 2024) uses an LLM to generate a video editing program pipeline based on the user instruction. In our work, we employ a VLM as a “VFX assistant” that, based on a short user instruction, provides a comprehensive description of the edited video along with the prominent objects present in the scene.
Professional Software For Video Animation In professional visual effects production, tools such as Autodesk Maya (Autodesk, INC., [n. d.]), Blender (Community, 2018), Unreal Engine (Epic Games, [n. d.]), Adobe After Effects (Christiansen, 2013) and Houdini (SideFX Houdini FX. SideFX, [n. d.]) are widely used for creating and compositing complex visual effects. These tools provide artists with precise control over object modeling, animation, and integration into video footage. While powerful, they require significant expertise, extensive manual intervention, and detailed inputs, such as 3D scene reconstruction or motion tracking. All the aforementioned software count on the user to provide the 3D assets. Even though creating input 3D assets has become an easier task to solve, using new datasets (Deitke et al., 2023; Qiu et al., 2024), or generating assets based on a user prompt (For example, Meshy.ai and Alpha3D), still, 3D physical elements (like fluid or explosion) or global/multi-object effects present a significant challenge. In this work, we take the first steps towards a novice user-friendly workflow.
Figure 3. DynVFX pipeline. Top row: Given an input video $\mathcal{V}_{\text{orig}}$, we apply DDIM inversion (see Sec. 3) and extract spatiotemporal keys and values $[\mathbf{K}_{\text{orig}},\mathbf{V}_{\text{orig}}]$ from the original noisy latents. Given the user instruction $\mathcal{P}_{\text{VFX}}$ we instruct the VLM to envision the augmented scene and output the text edit prompt $\mathcal{P}_{\text{comp}}$, prominent object descriptions $\mathcal{O}_{\text{orig}}$ that are used to mask out the extracted keys and values and target object descriptions $\mathcal{O}_{\text{edit}}$. Bottom row: We estimate a residual $\bm{x}_{\text{res}}$ to the original video latent ($\bm{x}_{\text{orig}}$). This is done by iteratively applying SDEdit with our Anchor Extended Attention, segmenting the target objects ($\mathcal{O}_{\text{edit}}$) from the clean result, and updating $\bm{x}_{\text{res}}$ accordingly.
@@ 3. Preliminaries
Diffusion probabilistic models (Ho et al., 2020; Sohl-Dickstein et al., 2015) are a class of generative models that aim to learn a mapping from noise $\bm{x}_{T}\sim\mathcal{N}(0,I)$ to a data distribution $q$. Starting from a Gaussian i.i.d noise $\bm{x}_{T}\sim\mathcal{N}(0,I)$, the diffusion model $\Phi$ is applied iteratively through a sequence of denoising steps, ultimately producing a clean output sample $\bm{x}_{0}$.
Recently, a new class of latent text-to-video (T2V) models, built on Diffusion Transformers (DiTs) (Peebles and Xie, 2023), has gained significant popularity. These models comprise multi-modal blocks (MMDiT (Esser et al., 2024)) that allow for joint processing of both text and image modalities, allowing each to inform and refine the other’s representations. To process both modalities together, first, a pre-trained encoder compresses an RGB video $\mathcal{V}$ both spatially and temporally to a latent space. Then, the latent is patchified, and the resulting tokens are concatenated to the text tokens produced by the text encoder (Raffel et al., 2023).
In each MMDiT block, text tokens and spatiotemporal tokens are projected into queries, keys and values using separate sets of weights for each modality, and the sequences of the two modalities are concatenated as a joint input for the attention operation: $\mathbf{Q}=[\mathbf{Q}_{\text{text}},\mathbf{Q}_{\text{spatio}}]$, $\mathbf{K}=[\mathbf{K}_{\text{text}},\mathbf{K}_{\text{spatio}}]$ and $\mathbf{V}=[\mathbf{V}_{\text{text}},\mathbf{V}_{\text{spatio}}]$. The attention (Vaswani et al., 2017) operation computes the affinities between the d-dimensional projections $\mathbf{Q},\mathbf{K}$ to yield the output of the layer:
(1) $\mathbf{A}\cdot\mathbf{V}\ \text{where}\ \mathbf{A}=\texttt{Attention}(\mathbf{Q},\mathbf{K})=\texttt{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\right)$
To capture inter-token relationships, Rotary Position Embeddings (RoPE) (Su et al., 2021) are applied to the input to the attention operation.

### v2

@@ Abstract.
We present a method for augmenting real-world videos with newly generated dynamic content. Given an input video and a simple user-provided text instruction describing the desired content, our method synthesizes dynamic objects or complex scene effects that naturally interact with the existing scene over time. The position, appearance, and motion of the new content are seamlessly integrated into the original footage while accounting for camera motion, occlusions, and interactions with other dynamic objects in the scene, resulting in a cohesive and realistic output video. We achieve this via a zero-shot, training-free framework that harnesses a pre-trained text-to-video diffusion transformer to synthesize the new content and a pre-trained vision-language model to envision the augmented scene in detail. Specifically, we introduce a novel inference-based method that manipulates features within the attention mechanism, enabling accurate localization and seamless integration of the new content while preserving the integrity of the original scene. Our method is fully automated, requiring only a simple user instruction. We demonstrate its effectiveness on a wide range of edits applied to real-world videos, encompassing diverse objects and scenarios involving both camera and object motion [^1] Code will be made publicly available.. Project page: https://dynvfx.github.io/
@@
submissionid: 2542journalyear: 2025conference: SIGGRAPH Asia 2025 Conference Papers; December 15–18, 2025; Hong Kong, Hong Kongbooktitle: SIGGRAPH Asia 2025 Conference Papers (SA Conference Papers ’25), December 15–18, 2025, Hong Kong, Hong Kongdoi: 10.1145/3757377.3764008isbn: 979-8-4007-2137-3/2025/12ccs: Computing methodologies Computer vision
Figure 1. DynVFX augments real-world videos with new dynamic content described by a simple user-provided text instruction (shown in the top row under each input video). The framework automatically infers where the synthesized content should appear, how it should move, and how it should harmonize at the pixel level with the scene, without requiring any additional user input. The key idea is to selectively extend the attention mechanism in a pre-trained text-to-video diffusion model, enforcing the generation to be content-aware of existing scene elements (anchors) from the original video. This allows the model to generate content that naturally interacts with the environment, producing complex and realistic video edits in a fully automated way (bottom row).
@@ 1. Introduction
Incorporating computer-generated imagery (CGI) into real-world footage has been a transformative capability in film production, enabling the creation of visual effects that would be difficult or impossible to achieve otherwise. For instance, the seamless integration of CGI characters, such as Gollum in The Lord of the Rings or T-Rex in Jurassic Park, has empowered filmmakers to blend fantastical elements with real-world environments, resulting in immersive storytelling. Professional video editing software tools provide artists with precise control over visual effects integration into video footage (Autodesk, 2024; Blender, 2024; Unreal Engine, 2024; Adobe, 2024). However, they require significant expertise, manual effort, pre-made assets, and financial resources. To disentangle these inspiring storytelling capabilities from professional and economic status, we propose a new creative task: free-vocabulary, text-driven integration of newly generated dynamic content into real-world videos. Specifically, given an input video along with a short user instruction describing the new content (e.g., “add a massive whale”), our goal is to synthesize new dynamic objects or complex scene effects that naturally interact with the existing scene across the entire video (Fig. 1). Importantly, our focus is to achieve this without additional user input, allowing anyone to insert dynamic content into videos just by describing it.
We approach this fully automated task by inferring where the synthesized content should appear, how it should move, and how it should harmonize at the pixel level with the scene without requiring any additional user input. This is achieved by harnessing dynamic generative priors learned by a pre-trained text-to-video diffusion transformer (DiT) model and comprehensive scene understanding learned by a vision-language model without any fine-tuning or additional training. This contrasts with previous methods that rely on additional user input, such as VFX assets (Hsu et al., 2025) or per-frame references to specify the placement of the new content over time. These methods struggle with integrating complex dynamic content (e.g., synthesizing multiple articulated objects or global effects, as shown in the dinosaurs and tsunami examples in Fig. 5).
@@
Our task poses several new fundamental challenges. First, the generation process of the fixed pre-trained DiT model must be content-aware such that the position, appearance, and motion of the synthesized dynamic content integrates naturally with the original scene. This entails synthesizing objects that respect occlusions, maintain appropriate relative size and perspective relative to the camera position, and realistically interact with other dynamic objects. All of this must be achieved while maintaining the integrity of the original video, ensuring that new content enhances the scene without compromising its authenticity.
Existing text-based methods focus on controlling diffusion generation to preserve original scene aspects. However, this imposes a trade-off between preserving aspects of the original scene and adding new content to the scene: either over-fitting or under-fitting the original video in terms of appearance, motion, positioning, or scale. Attention manipulation within a UNet-based architecture has been widely used for content preservation in video editing tasks. Extending attention across frames in these models allows for global appearance coherence. Yet, spatiotemporal information is preserved only implicitly. Hence, in the context of UNet-based models, this has been used mainly for appearance transfer. We revisit extended attention in the context of DiT-based text-to-video models, revealing new insights into their internal representation. Unlike UNet-based models, we observe that keys/values in DiTs locally encode corresponding video patches across space and time. In particular, applying the same positional embedding to both source and target keys/values during the extended attention operation enables alignment between the generated and original content at corresponding spatiotemporal locations.
We take advantage of this property to enforce content-aware generation for new content placement. Specifically, we steer the localization of the edit through Anchor Extended Attention: we incorporate a specific set of keys/values extracted from the original video as additional context to the DiT-based model during sampling, enabling the model to focus on existing scene elements essential for correctly placing and integrating new content naturally.
Despite accurate placement from Anchor Extended Attention, it does not guarantee pixel-level alignment with the original scene. To achieve better harmonization, we iteratively refine the edit by repeatedly sampling with Anchor Extended Attention while updating only the edited regions in each iteration.
To allow a fully automated framework, a vision-language model is utilized to interpret the interaction and reason about integration. We guide the model to: (1) translate the user’s instructions into detailed scene descriptions and (2) identify both the prominent visual elements of the existing scene and the new content to be introduced. These identified objects are used in our framework for the application of a text-based segmentation model to aid the localization and harmonization of the new content.
@@
We demonstrate the effectiveness of our approach on a variety of edits applied to real-world videos. Our method supports a wide range of scene augmentations across various scenarios while maintaining realistic interaction, occlusion, lighting, and camera motion. It is worth noting that our method utilizes a publicly available text-to-video model, which exhibits a significant gap in video generation quality compared to recent state-of-the-art video models. Nevertheless, we observe that within our problem formulation and objectives, we can distill surprisingly powerful generative capabilities from this model.
@@
  • We introduce a new task of integrating newly generated dynamic content into real-world videos without relying on the user to provide complex references for the effect.
  • We propose a tuning-free, zero-shot, fully automated method that enables harmonized content integration while maintaining high fidelity to the original scene.
  • We demonstrate state-of-the-art results, achieving the best trade-off between synthesizing new dynamic elements and maintaining high fidelity to the original content.
@@ 2. Related Work
Leveraging Text-to-Image Models. With the advancement of text-to-image (T2I) models, techniques for image manipulation based on such models have evolved rapidly. Among these advancements, notable progress has been made in the task of instruction-based image editing. Several works (Zhang et al., 2024d; Sheynin et al., 2024; Brooks et al., 2023; Zhang et al., 2023a) have proposed directly fine-tuning generative models on pairs of original and edited images coupled with user-provided instructions.
@@
Object insertion in images falls under the broader umbrella of instruction-based editing techniques. Recent object insertion methods (Canberk et al., 2024; Wasserman et al., 2025; Winter et al., 2024) leverage inpainting models to create paired image datasets, which are then used to fine-tune image editing models. However, extending these approaches to videos presents significant challenges. In particular, generating large-scale instruction-paired video datasets can be expensive in both time and computational resources, as it requires manual effort to annotate frames. This cost and complexity make it challenging to adapt existing image-based methodologies directly to the video domain.
Several works explore the manipulation of internal feature representation in T2I models for image and video editing tasks. In U-Net–based architectures, attention keys and values were shown to encode global appearance information (Cao et al., 2023), which several video editing methods (Qi et al., 2023; Wu et al., 2023; Khachatryan et al., 2023) exploit by extending self-attention across frames to enforce global appearance consistency. For localized appearance coherency, MasaCtrl (Cao et al., 2023) and Consistory (Tewel et al., 2024b) mask and selectively replace or extend attention keys and values. Add-It (Tewel et al., 2024a) employs global weighted extended attention for object insertion in images. FlowEdit (Kulikov et al., 2024) uses pre-trained flow models for inversion-free local edits. In contrast to the methods mentioned above that may fail to preserve the original content, we propose applying extended attention only to specific regions of the source scene, allowing generation to focus on essential elements.
Figure 2. Pipeline. Top (pre-processing). Given an input video $\mathcal{V}_{\text{orig}}$ and instruction $\mathcal{P}_{\text{VFX}}$, we (i) Apply our VLM protocol for instruction interpretation, to yield a comprehensive scene description $\mathcal{P}_{\text{comp}}$, original objects $\mathcal{O}_{\text{orig}}$ and target object $\mathcal{O}_{\text{edit}}$ descriptions. (ii) DDIM invert $\mathcal{V}_{\text{orig}}$ to extract spatiotemporal keys/values $\mathbf{K}_{\text{orig}},\mathbf{V}_{\text{orig}}]$. Bottom (editing). We initialize the composed latent with $\bm{x}_{\text{comp}}=\bm{x}_{\text{orig}}$ and iterate over a list of descending noise levels $t=\tilde{T}\!\rightarrow\!T_{\min}$ used for noising $\bm{x}_{\text{comp}}$. At each iteration $t$ we: (i) noise $\bm{x}_{\text{comp}}$ to noise level $t$, and sample with Anchored Extended Attention, to output $\bm{\hat{x}_{\text{comp}}}$. (ii) Update $\bm{x}_{\text{comp}}$ within the new contents masked regions $\bm{{M}_{\text{VFX}}}$ by adding the residual $\bm{x}_{\text{res}}=\bm{{M}_{\text{VFX}}\cdot(\hat{x}_{\text{comp}}-x_{\text{orig}}})$ to $\bm{x}_{\text{orig}}$. Repeating this loop gradually integrates the new content, yielding the edited video $\mathcal{V}_{\text{comp}}$.
@@
Controllable Video Generation and Editing. Recently, numerous methods have been developed to incorporate various forms of control signals into video generation pipelines. Several video-to-video methods propose to condition the generation on per-frame spatial maps such as depth maps and edge maps (Wang et al., 2023; Zhang et al., 2024b). A common approach in video editing is to enforce the preservation of the original scene layout and motion for various tasks. Several consistent video editing works (Bar-Tal et al., 2022; Lee et al., 2023; Geyer et al., 2023; Cong et al., 2023) restrict the preservation of scene structure and motion, and rely on a 2D generative prior. Therefore, they cannot support large structural deviations or motion adaptation. (Yatim et al., 2024; Zhao et al., 2024; Park et al., 2025; Jeong et al., 2024; Materzyńska et al., 2024) proposed to utilize a text-to-video (T2V) model for the task of motion transfer by controlling diffusion video generation to preserve original scene motion.
(Ku et al., 2024; Ouyang et al., 2024) adopt a two-stage pipeline: first, apply an off-the-shelf image editing method to the video’s first frame, then use an image-to-video generation model to propagate that edit through time via temporal feature injection. All of these methods fail to insert new dynamic content. Our approach focuses on integrating additional dynamic elements into the video.
Reference-Based Video Content Insertion. Recent methods (Bar-Tal et al., 2024; Ma et al., 2024; Zhang et al., 2024c; Zi et al., 2025; Zhou et al., 2023; Mou et al., 2024; Tu et al., 2025; Bian et al., 2025) have explored the adaptation of video models for video inpainting by conditioning on a masked video and a corresponding binary mask. This setup encourages the model to preserve unmasked information while generating new content in the masked region. Some methods, such as ReVideo (Mou et al., 2024) and VideoAnydoor (Tu et al., 2025), additionally guide the motion of the new content based on user-provided motion trajectories and bounding boxes.
VideoDoodles (Yu et al., 2023) combines hand-drawn animations with video footage in a scene-aware manner by tracking a user-provided planar canvas in 3D. However, it requires key-frame edit placement. Our task fundamentally differs from these methods, which all require explicit per-frame control over where (via per-frame masks (Bian et al., 2025) or bounding boxes (Mou et al., 2024)) or how the new content will move (via motion trajectories (Mou et al., 2024)). In contrast, our method does not constrain either aspect and requires only a simple text instruction as input. The locations and dynamics in our framework emerge from the generative prior of the T2V model. Furthermore, the reliance of these methods on user-provided per-frame annotations is impractical for integrating complex dynamics (e.g., Fig. 5 tsunami and dinosaurs). While a static object can be masked with a simple bounding box, manually defining per-frame masks for complex motion or interactions is extremely difficult.
Other approaches (KlingAI, 2024; Pika, 2024) reference the asset itself by requiring the user to additionally input an image of the object. This approach limits their applicability whenever a suitable example cannot be found for the given video. For instance, these methods struggle to generate global effects, making them less preferable than text-based insertion methods.
@@
Language Models for Video Content Creation. Advancements in Vision-Language Models (VLMs) have enabled methods to utilize such models in various video-related tasks. Some methods (Chen et al., 2024; Yang et al., 2024) use VLMs to produce detailed video captions from a series of frames, which are then utilized to train T2V generative models. Other methods utilize such models for achieving better generation controllability. For instance, VideoDirectorGPT (Lin et al., 2023) utilizes a VLM for multi-scene video generation by training diffusion adapters to incorporate additional conditioning inputs, while LVD (Lian et al., 2023) incorporates layout guidance from the VLM during the sampling process. AutoVFX (Hsu et al., 2025) uses an LLM to generate a video editing program pipeline based on the user’s instructions. In our work, we employ a VLM as a “VFX assistant”. Based on a short user instruction, it yields a comprehensive description of the edited video along with the prominent objects present in the scene.
@@ 3. Preliminaries
Diffusion probabilistic models (Ho et al., 2020; Sohl-Dickstein et al., 2015) are a class of generative models that aim to learn a mapping from noise $\bm{x}_{T}\sim\mathcal{N}(0,I)$ to a data distribution $q$. Starting from a Gaussian i.i.d. noise sample $\bm{x}_{T}\sim\mathcal{N}(0,I)$, the diffusion model $\Phi$ is applied iteratively through a sequence of denoising steps, ultimately producing a clean output sample $\bm{x}_{0}$.
Recently, a new class of latent T2V models, built on Diffusion Transformers (DiTs) (Peebles and Xie, 2023), has gained significant popularity, as DiTs enhance spatial coherence and enable training for arbitrary aspect ratios and video lengths. Unlike previous U-Net–based architectures, which utilize interleaving 2D spatial attention and 1D temporal attention layers, DiT treats the entire video as a long token sequence, where both text and spatiotemporal tokens are processed with full self-attention layers, a mechanism often referred to as 3D attention. In each DiT block, text tokens and spatiotemporal tokens are projected into queries, keys, and values using separate sets of weights for each modality, and the sequences of the two modalities are concatenated as a joint input for the attention operation. The attention (Vaswani et al., 2017) operation computes the affinities between the d-dimensional projections $\mathbf{Q}$ and $\mathbf{K}$ to yield the output of the layer:
(1) $\mathbf{A}\cdot\mathbf{V}\ \text{where}\ \mathbf{A}=\texttt{Attention}(\mathbf{Q},\mathbf{K})=\texttt{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\right)\,.$
To capture inter-token relationships, Rotary Position Embeddings (RoPE) (Su et al., 2024) are applied to the input queries and keys in the attention operation. In this work, we utilize a publicly available DiT-based model (Yang et al., 2024), CogVideoX, for augmenting real-world videos with newly generated dynamic content in a zero-shot manner.
Figure 3. Controlling Fidelity to the Original Scene Using Different Extended Attention Mechanisms. (a-b) SDEdit suffers from the original scene preservation/edit fidelity trade-off. (c-e) Three Extended Attention variants during sampling demonstrate different control levels: Full Extended Attention closely reconstructs the input scene, Masked Extended Attention proves too constrained in overlapping regions despite allowing new content emergence, and our Anchor Extended Attention. achieves optimal results by applying dropout – extending attention only at sparse points within selected regions.